let's assume a right triangle ABC. A is the angle theta but we don't know the degree. So a=5 that is the opposite side of theta, b=adjacent side that is unknown, and c=29 representing the hypothenus.
Determine the value of B. Round to two decimal places.

Answer by nerdybill(7384)

let's assume a right triangle ABC. A is the angle theta but we don't know the degree. So a=5 that is the opposite side of theta, b=adjacent side that is unknown, and c=29 representing the hypothenus.
Determine the value of B. Round to two decimal places.
.
Since you have a "right triangle", you can apply Pythagorean theorem:
a^2 + b^2 = c^2
5^2 + b^2 = 29^2
b^2 = 29^2 - 5^2
b^2 = 841 - 25
b^2 = 816
b = 28.57